OpenStudy (superhelp101):

The same amount of heat is added to 100 g of water and 1 kg of water starting at the same temperature. What will happen? The 100 g of water will reach a higher temperature. The 1 kg of water will reach a higher temperature. They will both reach the same temperature. There will be no significant change to either.

OpenStudy (superhelp101):

@Abhisar !

OpenStudy (abhisar):

What do you think it should be ?

OpenStudy (superhelp101):

ok i am think about The 100 g of water will reach a higher temperature.

OpenStudy (abhisar):

Why ?

OpenStudy (superhelp101):

because 100g of water will reach a higher temperature cause the formula heat= (mass h2o) (sp. ht. H2o) (delta T) if the 2 are equal the the delta T for the 100 g solution will have 10 x as large as delta for 1000g solution. sorry w-fi is bad it is taking forever.

OpenStudy (superhelp101):

i am new to this lesson so i not to sure

OpenStudy (abhisar):

\(\large \sf H=m \times C\times\delta T\), here C(specific heat) and H(energy) is constant so, -->\(\large\sf \delta T\propto \frac{1}{m}\)

OpenStudy (abhisar):

So, more the mass less will be the change in temperature and less the mass more will be the change in temperature.

OpenStudy (superhelp101):

oooooh! thank you!

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)